OpenStudy (anonymous):

can someone hlep me with this math problem? serie : an=(alpha*n)/(n+1), n>=1 (alpha is from R-real number). I must determine if the series is monotone and bouded thank you very much

OpenStudy (anonymous):

\[a _{n}=\frac{ \alpha*n }{ n+1 },n \ge1, \]\[, \alpha \in \mathbb{R}\]

OpenStudy (swag):

we would be glad to help but please post in the math group :)

OpenStudy (anonymous):

i m very sorry i didn t know how to find the groups section so i assumed that this is a general thread, thank you. i now noticed the url and found out that in am in the feedback group for the website, sorry again

OpenStudy (turingtest):

No worries, but I'm gonna close this Q since it's in the wrong group.

OpenStudy (anonymous):

yup, no problem, please do so, before that i want to ask you a stupid question where i can find the math group, it seems i can t figure it out, i am sory for the troubles created

OpenStudy (turingtest):

No apologies necessary, and welcome to OpenStudy! http://openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

Thank you very much